OpenStudy (anonymous):

Which ordered pairs in the form (x, y) are solutions to the equation (x + 7)y = 10? Choose all answers that are correct. A. (−9, −5) B. (−8, 10) C. (−2, 2) D. (3, 5)

OpenStudy (anonymous):

just plug in the numbers for x and y. a.(-9+7)-5=10? b.(-8+7)10=10? c.(-2+7)2=10? d.(3+7)5=10? now, calculate this and whichever one gives u 10 are correct